### v2

# Quantum-impurity sensing of altermagnetic order

Abstract
Quantum sensing with individual spin defects has emerged as a versatile platform to probe microscopic properties of condensed matter systems. Here we demonstrate that quantum relaxometry with nitrogen-vacancy (NV) centers in diamond can reveal the anisotropic spin dynamics of altermagnetic insulators together with their characteristic spin polarised bands. We show that the distance and orientation dependent relaxation rate of a nearby quantum impurity encodes signatures of momentum space anisotropy in the spin diffusion response, a hallmark of altermagnetic order. This directional sensitivity is unprecedented in the landscape of quantum materials sensing, and it enables the distinction of altermagnets from conventional antiferromagnets via local, noninvasive measurements. Our results could spark new NV-sensing experiments on spin transport and symmetry breaking in altermagnets, and highlight the role of NV orientation to probe anisotropic phenomena in condensed matter systems.
Introduction – Altermagnetism is a recently identified magnetic phase characterized by a unique spin arrangement of electrons in solids [1, 2, 3, 4]. Like antiferromagnets (AFMs), altermagnets (ALMs) exhibit vanishing net magnetization. However, unlike AFMs, ALMs exhibit time-reversal broken and spin-polarised bands, a characteristic of ferromagnets. Unlike the s-wave-like isotropic spin polarisation in ferromagnets, the spin polarisation in ALMs takes d, g, or i-wave form, i.e. a momentum-space structure of spin-split bands with energy iso-surfaces that are anisotropic, and intersect at 2, 4, or 6 spin-degenerate nodes [1]. Altermagnetic order was proposed in a wide range of materials [2, 5], including metals and insulators, and two dimensional systems [6], and it has been experimental confirmed in a few materials [7, 8, 9, 10] and in some cases remains stable under ambient conditions [9, 11, 12]. Its discovery motivated by predicition and observation of unconventional crystal anomalous Hall effect [13, 14, 15, 16] opens up exciting research avenues [1, 2, 5, 3], including magnetoresistance effects [17], topological [6, 18], multiferroic properties [19, 20, 21], and interplay with superconductivity [22], to name a few.
In this Letter, we propose a non-invasive sensing protocol based on the relaxometry properties of quantum spin impurities, such as nitrogen-vacancy (NV) centers in diamond [23, 24, 25, 26, 27, 28], in order to distinguish altermagnets (ALMs) from conventional magnetic materials. Specifically, the dependence of the relaxation rate with the relative orientation between the NV principal axis and the Néel vector of the material changes with the distance between the NV and the sample – a feature not present in conventional magnets [29, 30, 31]. The spin diffusion coefficients of ALM insulators can then be obtained by analyzing the decay of the impurity’s relaxation rate at different orientations with the distance from the sample material, as schematized in Fig. 1(a). We propose a contrast function, exemplified in Fig. 1(b), that depends exclusively on the diffusion coefficients, and exhibits an excursion of up to 27 % with the NV-sample distance, for commercially available NVs, whilte it would be completely flat for conventional magnets, such as AFMs and ferromagnets. Our scheme is a feasible and non-invasive tool for probing spin diffusion in ALM insulators, a task so-far not accomplished in experiments, and which is valuable for devising new applications in which spin transport is pivotal, such as spin-based information processing.
Figure 1: (a) Relaxometry of spin diffusion in altermagnets. The relaxation rate $\Gamma$ of a quantum impurity with transition frequency $\omega$ is monitored for different relative orientations, described by the set of angles $\{\Phi\}=(\theta,\varphi)$, and at different distances $d$ to the altermagnet. The spin diffusion length $l_{0}$ sets the relevant distance scale. (b) The relaxation contrast over a set of relative orientations exhibits a non-trivial dependence with the distance, which is a fingerprint unique to altermagnetism. Its value can increase up to $\sim 27$ %.
Quantum impurity relaxometry – We consider the setup of Fig. 1. A quantum impurity (QI), such as a NV center, is placed in the proximity of a two-dimensional magnetic material. In the following, we will keep our derivation general, and specialize to NVs whenever necessary. The relaxation rate $T_{1}$ of the impurity is then composed by an intrinsic contribution (e.g. due to decay into phonons for NVs), and a contribution due to the noisy stray magnetic field of the sample. The QI magnetic moment, associated to a transition between two spin levels, couples to the stray magnetic field generated by the magnetic excitations of the material via dipole coupling. The noisy stray magnetic field causes transitions between the two levels with relaxation rate [29, 32]
$\Gamma[\omega]=\gamma^{2}\int dte^{-i\omega t}\langle\{\hat{B}^{(+)}_{\rm{QI}}(t),\hat{B}^{(-)}_{\rm{QI}}(0)\}\rangle/2,$ (1)
where $\{\cdot,\cdot\}$ is the anticommutator, $\hat{B}^{(+)}_{\rm{QI}}(t)=\hat{B}_{x,\rm{QI}}+\hat{B}_{y,\rm{QI}}$ is the magnetic field operator associated with the spin fluctuations of the material, $\hbar\omega$ the energy level splitting of the quantum impurity and $\gamma$ the gyromagnetic ratio of the dipole transition. By measuring $T_{1}$, it is then possible to infer the additional decay of the NV due to the magnetic noise.
The stray magnetic field correlations carry information about single and multiple magnon scattering processes in the sensed magnet [33, 34, 29, 35, 32, 36, 31]. For a NV with frequency $\omega$ lying within the magnon gap, single magnon processes will be off resonant and thus will not contribute to the relaxation rate. In this case, the main contribution to $\Gamma$ is the longitudinal correlation function $\langle\{\hat{s}_{\parallel}(\vec{r},t),\hat{s}_{\parallel}(\vec{r}^{\prime},0)\}\rangle$, where $\parallel$ refers to the component of the spin density operator $\hat{\vec{s}}$ parallel to the Néel vector. This correlation function is associated with two magnon processes and, thus, with spin transport. These conditions hold for a NV-vacancy center with $\omega\sim$ GHz [37], and an ALM in a weak bias field and with easy-axis anisotropy [38, 39, 40, 41].
Using the fluctuation-dissipation theorem [42], we can rewrite the relaxation rate (1) in terms of the imaginary part of the spin susceptibility $\chi^{\prime\prime}_{\parallel}(\omega,\vec{k})$ [34, 32]:
$\Gamma[\omega]=\frac{\hbar\gamma^{2}\tilde{\gamma}^{2}}{2}{\rm{coth}}\left(\frac{\hbar\beta\omega}{2}\right)\int\frac{d^{2}\vec{k}}{(2\pi)^{2}}\mathcal{C}_{\{\Phi\}}(d,\vec{k})\chi^{\prime\prime}_{\parallel}(\omega,\vec{k}),$ (2)
where $\tilde{\gamma}$ is the magnet’s gyromagnetic ratio and $d$ is the QI-sample distance. The geometric factor $\mathcal{C}_{\{\Phi\}}(d,\vec{k})=(2\pi)^{2}k^{2}e^{-2kd}\mathcal{F}_{\{\Phi\}}(\phi_{k})$ (with ${\rm{tan}}(\phi_{k})=k_{y}/k_{x}$) is obtained from the magnetostatic Green’s function of a dipole [43], which is commonly encountered in NV-based noise spectroscopy [34, 32, 36], and depends on the relative orientation between the NV principal axis (assumed to be in the $yz$ plane) and the in-plane Néel vector of the ALM via the function $\mathcal{F}_{\{\Phi\}}(\phi_{k})$, where the set of angles $\{\Phi\}=(\theta,\phi)$ describe the orientation of the NV principal axis and the Néel vector, as illustrated in Fig. 1. We remark that our formalism would also be suitable for Néel vectors with out-of-plane components. The coefficient $k^{2}e^{-2kd}$ in $\mathcal{C}_{\{\Phi\}}(d,\vec{k})$ introduces a momentum filter function: for $k\gtrsim 1/d$ the integrand is suppressed by the exponential factor $e^{-2kd}$, and at the same time it dampens it for $k\lesssim 1/d$ due to the term proportional to $k^{2}$. This basically means that we can single out the response function Eq. (2) around $k\simeq 1/d\equiv k_{d}$ in the momentum integral involved in the calculation of the relaxation rate (2), or in physical terms, only magnons with wavelengths $\simeq 1/d$ contribute. The central result of this Letter is that the relaxation rate in Eq. (2) can distinguish between spin diffusion in ALMs and in more conventional magnets such as antiferromagnets (AFMs), as a result of the anisotropy of the diffusion response function $\chi_{\parallel}(\omega,\vec{k})$ in $k$-space – a consequence of the alternating magnon band structure of an ALM [38, 40].
Spin diffusion in ALMs – We obtain the response function $\chi_{\parallel}$ in the diffusive limit following the phenomenological transport approach in [44, 45]. ALMs can be modeled through two magnetic sub-lattices, and thus host two species of magnons, indicated by $\alpha$ and $\beta$, which carry opposite spins [45]. When the system is perturbed away from thermal equilibrium, for example, due to a temperature gradient, or a magnetic field [46], magnons can diffuse, thus transporting spin. For each of the two ALM magnon species, we assume that the excess magnon density with respect to its equilibrium value, which we denoted by $\delta n_{\xi=\alpha,\beta}$, behaves diffusively:
$\partial_{t}\delta n_{\xi}+\nabla\cdot\vec{j}_{\xi}$ $=-\frac{1}{\tau_{s}}(\delta n_{\xi}\pm\chi_{0}h),$ (3)
where $h$ is a generalized force thermodynamically conjugated to $\delta n_{\xi}$, $\tau_{s}$ is the relaxation time, and $\vec{j}_{\xi}=-\overleftrightarrow{D}_{\xi}\cdot\nabla(\delta n_{\xi}\pm\chi_{0}h)$ is the magnon density current, which is given in terms of the diffusion tensor $\overleftrightarrow{D}_{\xi}$. The static spin susceptibility is indicated by $\chi_{0}$, and the associated chemical potential is $(\delta n_{\xi}/\chi_{0}\pm h)$ [47], with the different signs corresponding to the different magnon species [46]. Relying on a Boltzmann distribution for the steady-state magnon occupation [45], we can obtain the elements of the diffusion tensor [48], which depend on the magnon band structure, and the magnon momentum relaxation rate [41].
As a simple way to model magnon bands in a 2D ALM [49, 50, 51], we resort to the Lieb lattice of Fig. 2(a) with the Heisenberg Hamiltonian
$\hat{\mathcal{H}}$ $=J_{1}\sum_{\langle\vec{r},\vec{r}^{\prime}\rangle}\hat{\vec{S}}_{i}\cdot\hat{\vec{S}}_{j}+\sum_{\langle\langle\vec{r},\vec{r}^{\prime}\rangle\rangle_{\pm}}(J_{2}\pm\Delta)\hat{\vec{S}}_{i}\cdot\hat{\vec{S}}_{j}$ (4)
$+\frac{J_{\rm{EA}}}{2}\sum_{\vec{r}}\left(\hat{S}_{z,j}\right)^{2},$
where $J_{1}$ is the nearest-neighbor exchange between sublattices with opposite spins, $(J_{2}\pm\Delta)$ are the next-nearest-neighbor exchange, which depend on the direction, and $J_{\rm{EA}}$ is the on site easy-axis anisotropy. We then perform a Holstein-Primakoff approximation to first order [45, 48], and obtain a quadratic bosonic Hamiltonian describing the excitation on top of the Néel order. A standard diagonalization procedure yield the magnon frequencies $\omega_{\alpha,\beta}$ shown in Fig. 2(b) around the indicated path in the Brillouin zone. These magnon bands display the alternating band splitting which, in this model, is parametrized by $\Delta$. In order to measure spin diffusion, the QI transition frequency $\omega$ has to lie within the magnon gap $\Delta_{\rm{mag}}=4s(16J_{1}J_{\rm{EA}}+J_{\rm{EA}}^{2})^{1/2}/\hbar$, rendering any single-magnon process (e.g. annihilation of a magnon followed by a transition in the QI) off-resonant. This guarantees that longitudinal spin-correlations constitute the major contribution for Eq. (1). For a QI with $\omega\sim$ GHz and an ALM with exchange constant $J_{1}\sim 10$ meV, even a weak anisotropy $J_{\rm{EA}}$ would satisfy such requirement.
The spin model considered here illustrates a d-wave ALM, with the magnetic lattice exhibiting a 4-fold rotation symmetry [39], which is also present in the magnonic band structure [52]. In turn, we can show that the magnon diffusion tensor has the general form $D_{\alpha,xx}=D_{\beta,xx}=D_{\alpha,zz}=D_{\beta,zz}=D_{1}$, while $D_{\alpha,xz}=-D_{\beta,xz}=D_{2}$ [48]. Such relations should hold for any d-wave ALM, and are a direct consequence of the symmetries of the Heisenberg Hamiltonian. We furthermore notice that for $\Delta=0$, i.e., in the absence of the ALM band splitting, $D_{2}=0$, in which case the model describes a Liebe lattice AFM. The isotropic diffusion of tensor of such an ALM is a consequence of the Heisenberg Hamiltonian – the exchange constants are the same for all components of the spins, and are isotropic.
Figure 2: (a) Lieb lattice altermagnet. The different exchange constant are: $J_{1}$, the nearest neighbor exchange, and $J_{2}\pm\Delta$ the next-nearest neighbor exchanges. The system is assumed in an in-plane Néel state due to an in-plane easy-axis anisotropy $J_{\rm{EA}}$. (b) Magnon frequencies $\omega_{\alpha,\beta}$ along the indicated path in the Brillouin zone for $J_{2}=0.5J_{1}$, $\Delta=0.1J_{1}$, and $J_{\rm{EA}}=0.01J_{1}$.
The spin diffusion response function is then obtained from the set of coupled linear diffusion equations for $\mathcal{S}_{\parallel}=\hbar(-\delta n_{\alpha}+\delta n_{\beta})$, related to the longitudinal spin density, and $\mathcal{S}_{\perp}=\hbar(\delta n_{\alpha}+\delta n_{\beta})$, related to the total magnon density. The dynamics of the magnon densities are given by Eq. (3), and the diffusion tensor couples $\mathcal{S}_{\parallel}$ to $\mathcal{S}_{\perp}$. These equations can be solved in Fourier domain through $\mathcal{S}_{\parallel}(\omega,\vec{k})=\chi_{\parallel}(\omega,\vec{k})h$, with the response function is given by [48]
$\chi_{\parallel}(\omega,\vec{k})=\frac{\hbar\chi_{0}\left[D_{1}k^{2}+\frac{1}{\tau_{s}}-\frac{D_{2}^{2}k^{4}\cos^{2}(2\phi_{k})}{-i\omega+D_{1}k^{2}+1/\tau_{s}}\right]}{-i\omega+D_{1}k^{2}+\frac{1}{\tau_{s}}-\frac{D_{2}^{2}k^{4}\cos^{2}(2\phi_{k})}{-i\omega+D_{1}k^{2}+1/\tau_{s}}}.$ (5)
The dependence of $\chi_{\parallel}$ on the directionality in $k$-space and the factor $\propto k^{4}$ are traits unique of ALMs, not present in usual magnets [29, 30, 31].
Relaxation rate and contrast – The response function $\chi_{\parallel}$ in Eq. (5) is anisotropic in $k$-space and includes the dependence $k^{4}$. As a consequence, the dependence of the relaxation rate on the NV-sample distance $d$ and on the relative orientation $\{\Phi\}$ for ALMs is different than that of conventional magnets. To better highlight this feature, we find it useful to define the relaxation contrast
$\mathcal{C}[d]=\frac{\Gamma_{\rm{max}}[d]-\Gamma_{\rm{min}}[d]}{\Gamma_{\rm{max}}[d]+\Gamma_{\rm{min}}[d]},$ (6)
which depends on the maximum and the minimum of the relaxation rate over a set of relative orientations at a given QI-sample distance. For an AFM, the spin diffusion susceptibility is isotropic, and, as a result of that Eq. (6) will be independent of the distance $d$, while ALM exhibit a non-trivial dependence of the contrast with the QI-sample distance. To further analyze the contrast and the QI relaxation, we take the DC limit ($\omega\to 0$) [^1] This approximation requires that $\omega\ll D_{1}l_{0}/d^{2}$. We nevertheless emphasize that the plots shown in the paper are for a finite $\omega=1/\tau_{s}$., which should be valid for a NV operating in the GHz regime.
The contrast is maximized when the QI axis is aligned with the magnet’s plane, i.e. $\theta=\pi/2$, cf. Fig. 1. We then consider variations in $\varphi$. The maximum relaxation rate occurs for $\varphi=0$, while the minimum is at $\varphi=\pi/2$. The contrast for $d\gg\sqrt{D_{1}\tau_{s}}$ is then given by $\mathcal{C}(d)|_{d\gg l_{0}}=11/21\approx 0.52$, which is independent of any microscopic parameter. In fact, the relaxation rate in this limit is
$\Gamma[\omega]|_{d\gg l_{0}}\propto\frac{l_{0}^{4}}{d^{4}}\int_{0}^{2\pi}d\phi_{k}\mathcal{F}_{\theta;\vartheta,\varphi}(\phi_{k}),$ (7)
where we have defined $l_{0}=\sqrt{D_{1}\tau_{s}}$, the isotropic spin diffusion length. Such a relaxation rate has exactly the same form one would get from a magnet with isotropic magnon bands [29, 30]. The value of the contrast is independent on whether the material is an ALM or an AFM, a consequence of the anisotropy of the magnetic noise driving the QI with respect to the orientation angle $\varphi$.. The Néel vector sets the direction of the longitudinal modes probed by relaxometry, rendering the relaxation rate intrinsically anisotropic. ALMs induces further anisotropy contributions due to the magnon splitted bands, which, according to Eq. (5), is washed away for wave-vectors $k\ll 1/l_{0}$. Thus, the NV is insensitive to such additonal trait intrinsic to ALMs when placed at distances $d\gg l_{0}$.
At close distances $d\ll l_{0}$, the QI becomes sensitive to the intrinsic anisotropy of spin diffusion in ALMs, since the relaxation rate has a different dependence on the orientation angles. In this limit the relaxation rate is given by  [^2] The singularity of the integrand at $\phi_{k}=0,\pi/2,$... and $D_{2}=D_{1}$ is an artifact of the approximation $\omega\sim 0$ that was taken to write Eq. (8). The frequency $\omega$ is never zero, and thus the integrand is always finite. All plots shown in the paper are generated with the full expression including a non-vanishing frequency, which is shown in the SI.
$\Gamma[\omega]|_{d\ll l_{0}}\propto\frac{D_{1}^{2}l_{0}^{2}}{d^{2}}\int_{0}^{2\pi}d\phi_{k}\frac{\mathcal{F}_{\theta;\vartheta,\varphi}(\phi_{k})}{D_{1}^{2}-D_{2}^{2}\cos^{2}(2\phi_{k})}.$ (8)
The anisotropy of the response function Eq. (5) induces an additional dependence of the integrand from directionality in $k$-space in the above equation. In turn, the contrast in this limit is given by
$\mathcal{C}(d\ll l_{0})=\frac{5D_{1}(D_{1}-\sqrt{D_{1}^{2}-D_{2}^{2}})+3D_{2}^{2}}{3D_{1}(D_{1}-\sqrt{D_{1}^{2}-D_{2}^{2}})+9D_{2}^{2}}.$ (9)
In the case of an AFM, i.e. by setting $D_{2}=0$, the contrast yields $11/21$, which is the same value it has at distances $d\gg l_{0}$. For an ALM, the contrast behaves non-trivially with distance, exhibiting a larger value at closer distances. For $D_{2}\approx D_{1}$ (its maximum value), the contrast at closer distances is $8/12\approx 0.67$, which is an increase of 2727% compared to the large distance limit. This is the case shown in Fig. 1(b), for which we took $D_{2}\approx 1$. In general, the contrast depends on the ratios $D_{2}/D_{1}$ and $d/l_{0}$, and on the product $\omega\tau_{s}$ [48], thus its behavior with respect to the other parameters can also be exploited for extracting intrinsic diffusion parameters. We would like to emphasize that the behavior of the contrast showed here is a consequence of the anisotropy of the response function in Eq.  (5), generated by the off-diagonal component of the diffusion tensor $D_{2}$.
We show in Fig. 3 the relaxation rate $\Gamma$ and the contrast $\mathcal{C}$ in as a function of the distance $d$ for different values of the parameter $\Delta$. The relaxation rate is shown in 3(a) in units of the characteristic rate $\Gamma_{c}=\hbar\chi_{0}\gamma^{2}\tilde{\gamma}^{2}k_{b}T\tau_{s}/l_{0}^{4}$. The relaxation rate for an ALM, depicted by the dark curves, exhibits a different curvature at the cross-over region $d=l_{0}$ when compared to an usual AFM, depicted in light blue, due to the term $\propto k^{4}$ appearing in the susceptibility Eq. (5). Furthermore, as depicted in Fig. 3(b), the contrast increases as the distance $d$ is decreased, a trait that is unique to ALMs and is a consequence of the anisotropy of the spin diffusion susceptibility.
Figure 3: (a) QI relaxation rate as a function of the distance to the sample $d$ in units $\Gamma_{\rm{c}}$, given in the text, for different relative orientations. (b) Relaxation contrast defined in Eq. (6) as a function of the distance $d$ for different values of the parameter $\Delta$. Parameters are: $J_{2}=-0.8J_{1}$, $J_{\rm{EA}}=0.01J_{1}$, and $\omega\tau_{s}=1$.
[…]
Perspectives - In this Letter, we discussed the quantum noise spectroscopy of insulating ALMs using NV centers, where the carriers of spin in the material are localized. In a similar way, quantum noise spectroscopy can also be used to study the properties of metallic ALMs (MALMs). In this case, additional considerations arise. Similar to the noise spectroscopy of generic metallic systems, one must account for the Johnson noise due to current fluctuations [33, 34, 62]. The notable feature of MALMs, however, is the mobile nature of magnetic moments in conjunction with strong spin-orbit interaction, leading to a qualitatively distinct dynamical spin response function and the resulting magnetic noise. The presence of spin-momentum locking leads to a rich dependence of the relaxation on the orientation of the NV center. Moreover, nodal structure of altermagneto electronic structure [39], significantly alters the magnetic noise spectrum. The gapless nature of the magnetic excitations also makes $T_{2}$-based spectroscopy a particularly powerful tool for probing the spin dynamics in such systems, as discussed below.
An alternative approach to probing ALMs is through $T_{2}$ relaxometry [63]. The typically stronger $T_{2}$ signal, as compared to $T_{1}$ noise, together with its operational low-frequency range, makes it well-suited for studying systems with low-energy and gapless modes. These include, the longitudinal diffusive modes, hydrodynamic sound modes, as well as nodal magnons in MALMs mentioned above. Furthermore, recent advances in NV-based sensing allow for multi-qubit noise spectroscopy, which can provide spatially resolved information on magnetic noise correlations [64, 28]. This technique offers powerful means to probe non-invasively spatio-temporal spreading of correlations in ALMs [65].
Acknowledgments - H.H. and J.M. are grateful to E. Demler, P. Dolgirev, S. Gopalakrishnan for related discussions on NV sensing. VASVB acknowledges financial support from the Contrat Triennal 2021-2023 Strasbourg Capitale Europeenne. JS and LS acknowledge support by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - DFG (Project 452301518) and TRR 288 – 422213477 (project A09 and A12). LS acknowledges support by the ERC Starting Grant No. 101165122. JM acknowledges support by the DFG through the grant HADEQUAM-MA7003/3-1. We acknowledge support from the Dynamics and Topology Centre, funded by the State of Rhineland Palatinate.
Supplementary Material: Quantum impurity sensing of altermagnets
I Magnetostatic Green’s function and the geometric factor of the relaxation rate
In this section we present explicitly the relation between the stray magnetic field $\hat{\vec{B}}_{\rm{QI}}$ in Eq.(1) and the spin density $\hat{\vec{s}}$ and obtain explicitly the geometric factor $\mathcal{C}_{\{\Phi\}}(d,\vec{k})$ of Eq. (2). The stray magnetic field $\hat{\vec{B}}_{\rm{QI}}$ is given by the the stray magnetic field $\hat{\vec{B}}$ written in the quantum impurity (QI) principal axis frame. We keep to the setup of Fig. 1, and consider that the QI principal axis lies in the yz plane, such that $\hat{\vec{B}}_{\rm{QI}}=\mathcal{R}_{x}(\theta)\cdot\hat{\vec{B}}$, where $\mathcal{R}_{x}(\theta)$ is a rotation around the $x$ axis.
The field $\hat{\vec{B}}(\vec{r},t)$ and the spin density $\vec{s}$ are related by the magnetostatic Green’s function [26]:
$\hat{\vec{B}}(\vec{r},t)=\tilde{\gamma}\int d^{3}\vec{r}^{\prime}\overleftrightarrow{\mathcal{D}}(\vec{r},\vec{r}^{\prime})\cdot\hat{\vec{s}}(\vec{r}^{\prime},t),$ (S1)
where $\overleftrightarrow{\mathcal{D}}(\vec{r},\vec{r}^{\prime})=-\nabla_{\vec{r}}\nabla_{\vec{r}^{\prime}}(1/|\vec{r}-\vec{r}^{\prime}|)$. The elements of the magnetostatic Green’s function for a thin film read
$\mathcal{D}_{jl}(\vec{r},\vec{r}^{\prime})$ $=-\frac{1}{2\pi}\int d^{2}\vec{k}\frac{k_{j}k_{l}}{k}e^{-k|y-y^{\prime}|}e^{i\vec{k}\cdot(\vec{\rho}-\vec{\rho}^{\prime})}\quad\quad(j,l=x,z),$ (S2)
$\mathcal{D}_{yj}(\vec{r},\vec{r}^{\prime})$ $=-\frac{i\rm{sign}(y-y^{\prime})}{2\pi}\int d^{2}\vec{k}k_{j}e^{-k|y-y^{\prime}|}e^{i\vec{k}\cdot(\vec{\rho}-\vec{\rho}^{\prime})}\quad\quad(j=x,z),$
$\mathcal{D}_{zz}(\vec{r},\vec{r}^{\prime})$ $=-\frac{1}{2\pi}\int d^{2}\vec{k}ke^{-k|y-y^{\prime}|}e^{i\vec{k}\cdot(\vec{\rho}-\vec{\rho}^{\prime})},$
where $k=\sqrt{k_{x}^{2}+k_{z}^{2}}$ and $\vec{\rho}=x\vec{e}_{x}+z\vec{e}_{z}$ (with a similar definition for $\vec{\rho}^{\prime}$). Furthermore, the spin density $\hat{\vec{s}}(\vec{r}^{\prime},t)$ is proportional to $\delta(y)$, and can be written as $\hat{\vec{s}}(\vec{r},t)=\mathcal{R}_{y}(\varphi)\hat{\vec{s}}_{N}(\vec{r},t)$, where $\hat{\vec{s}}_{N}$ is the spin density written in a frame of the Néel vector, and $\mathcal{R}_{y}(\varphi)$ is a rotation around the $y$ axis. We can then write $\hat{B}_{{\rm{QI}}}^{(-)}=\hat{B}_{{\rm{QI}},x}-i\hat{B}_{{\rm{QI}},y}$, which is required for calculating the relaxation rate, as
$\hat{B}_{{\rm{QI}}}^{(-)}(d,t)$ $=\tilde{\gamma}\int d^{2}\vec{\rho}^{\prime}\Big{[}\mathcal{A}^{-}_{\theta,\varphi}(\vec{\rho}^{\prime},d)\hat{s}^{(-)}(\vec{\rho}^{\prime},t)+\mathcal{B}^{-}_{\theta,\varphi}(\vec{\rho}^{\prime},d)\hat{s}^{(+)}(\vec{\rho}^{\prime},t)+\mathcal{C}^{-}_{\theta,\varphi}(\vec{\rho}^{\prime},d)\hat{s}_{\parallel}(\vec{\rho}^{\prime},t)\Big{]},$ (S3)
where the components of the spin operators are in the frame of the Néel vector. The relaxation rate in Eq. (1) has contribution from different correlators. As argued in the main text, we the most relevant for the setup we envision are the ones proportional to $\langle\{\hat{s}_{\parallel}(\vec{\rho}^{\prime},t),\hat{s}_{\parallel}(\vec{\rho}^{\prime,\prime},0)\}\rangle$, where $\{\cdot,\cdot\}$ denotes the anticommutator. We substitute Eq. (S3) into Eq. (1) and write the corresponding integrals in the Fourier domain keeping only the aforementioned spin correlators, which gives
$\Gamma[\omega]=\gamma^{2}\tilde{\gamma}^{2}\int\frac{d^{2}\vec{k}}{(2\pi)^{2}}\mathcal{C}_{\{\Phi\}}(d,\vec{k})C_{\parallel}(\vec{k},\omega),$ (S4)
where $C_{\parallel}(\vec{k},\omega)$ is the Fourier transform of the spin correlator. $\mathcal{C}_{\{\Phi\}}(d,\vec{k})$ can be obtained via the factor $\mathcal{C}^{-}_{\theta,\varphi}$ of Eq: (S3), which uses the Fourier domain form of the elements of the magnetostatic Green’s tensor. After some algebra we have
$\mathcal{C}_{\{\Phi\}}(d,\vec{k})$ $=(2\pi)^{2}k^{2}e^{-2kd}\mathcal{F}_{\{\Phi\}}(\phi_{k})$ (S5)
[…]
$\quad+2\sin(\varphi)\cos(\varphi)\sin\left(\phi_{k}\right)\cos\left(\phi_{k}\right)\left(\cos\left(\phi_{k}\right)-\sin(\theta)\right)\left(\sin(\theta)+2\cos\left(\phi_{k}\right)\right)$
The final form of the relaxation rate appearing in Eq. (2) is then obtained by relating $C_{\parallel}(\vec{k},\omega)$ to the imaginary part of the response function $\chi_{\parallel}$ via the fluctuation-dissipation theorem [42].
II Magnon diffusion tensor
We present here the full formula for the magnon diffusion tensor $\overleftrightarrow{D}$. We follow the procedured used for ferro- and antiferromagnets [44, 45]. Our starting point is the excess density $\delta n_{\xi}(\vec{r})$ of each magnon species $\xi=\alpha,\beta$ is
$\delta n_{\xi}(\vec{r})=\frac{1}{(2\pi)^{2}}\int d^{2}k\left(n_{\xi,\vec{k}}(\vec{r})-n_{\xi,k}^{0}\right),$ (S6)
where $n_{\xi,k}^{0}$ is the equilibrium distribution of a given magnon mode (the Bose-Einstein distribution). We can also write the excess magnon current is
$\vec{j}_{\xi,\rm{Tot}}=\frac{1}{(2\pi)^{3}}\int d^{2}k\vec{v}_{\xi\vec{k}}\left[n_{\xi,\vec{k}}(\vec{r})-n_{\xi,k}^{0}\right],$ (S7)
$\vec{v}_{\xi\vec{k}}=\partial\omega_{\xi\vec{k}}/\partial\vec{k}$ is the magnon velocity. The steady-state excess distribution is given then by the solution of the Boltzmann equation
$n_{\xi,\vec{k}}(\vec{r})-n_{\xi,k}^{0}=-\tau_{\xi,\vec{k}}\vec{v}_{\xi\vec{k}}\cdot\nabla n_{\xi,\vec{k}}(\vec{r}),$ (S8)
where $\tau_{\xi,\vec{k}}$ is the momentum relaxation rate.
The excess magnon current density can then be decomposed in two components [44]: one related to changes on the equilibrium magnon distribution $n_{\xi,k}^{0}$ due to the change of some parameter, such as the temperature, and the contribution due to spatial accumulation of magnons. To describe the magnon diffusive properties, we will focus on the latter which is given by
$\vec{j}_{\xi}=-\frac{1}{(2\pi)^{2}}\int d^{2}k\tau_{\xi,\vec{k}}\vec{v}_{\xi\vec{k}}\left[\vec{v}_{\xi\vec{k}}\cdot\nabla(n_{\xi,\vec{k}}(\vec{r})-n_{\xi,k}^{0})\right].$ (S9)
We can then write such a current in terms of a magnon diffusion tensor. For that, we consider the following form for the excess magnon distribution
$\delta n_{\xi,\vec{k}}(\vec{r})=n_{\xi,\vec{k}}(\vec{r})-n_{\xi,k}^{0}=n_{\xi,k}^{0}\varepsilon_{\xi,\vec{k}}g(\vec{r}),$ (S10)
where $\varepsilon_{\xi,\vec{k}}=\hbar\omega_{\xi,\vec{k}}$, and thus the total magnon density accumulation is
$\delta n_{\xi}(\vec{r})$ $=\frac{1}{(2\pi)^{2}}\int d^{2}k\left(n_{\xi,\vec{k}}(\vec{r})-n_{\xi,k}^{0}\right)$ (S11)
$=\frac{1}{(2\pi)^{3}}\int d^{2}kn_{\xi,\vec{k}}^{0}\varepsilon_{\xi,\vec{k}}g(\vec{r})=\frac{g(\vec{r})}{(2\pi)^{2}}\int d^{3}kn_{\xi,k}^{0}\varepsilon_{\xi,\vec{k}}.$
We call from now on
$I_{0}=\frac{1}{(2\pi)^{2}}\int d^{2}kn_{\xi,\vec{k}}^{0}\varepsilon_{\xi,\vec{k}},$ (S12)
and thus
$g(\vec{r})=\frac{\delta n_{\xi}(\vec{r})}{I_{0}}.$ (S13)
We have then
$\nabla\delta n_{\xi,\vec{k}}(\vec{r})=n_{\xi,k}^{0}\varepsilon_{\xi,\vec{k}}\nabla g(\vec{r})=\frac{n_{\xi,k}^{0}\varepsilon_{\xi,\vec{k}}}{I_{0}}\nabla\delta n_{\xi}(\vec{r}),$ (S14)
and thus Eq. (S9) reads
$\vec{j}_{\xi}=-\frac{1}{(2\pi)^{2}}\int d^{2}k\frac{n_{\xi,k}^{0}\varepsilon_{\xi,\vec{k}}}{I_{0}}\tau_{\xi,\vec{k}}\vec{v}_{\xi\vec{k}}\left[\vec{v}_{\xi\vec{k}}\cdot\nabla\delta n_{\xi}(\vec{r})\right]=-\frac{1}{I_{0}(2\pi)^{2}}\int d^{2}kn_{\xi,k}^{0}\varepsilon_{\xi,\vec{k}}\tau_{\xi,\vec{k}}\vec{v}_{\xi\vec{k}}\left[\vec{v}_{\xi\vec{k}}\cdot\nabla\delta n_{\xi}(\vec{r})\right].$ (S15)
We write then for each component of the magnon excess current
$j_{i,\xi}=-\left[\sum_{j}\frac{1}{I_{0}(2\pi)^{2}}\int d^{2}kn_{\xi,k}^{0}\varepsilon_{\xi,\vec{k}}\tau_{\xi,\vec{k}}v_{i,\xi\vec{k}}v_{j,\xi\vec{k}}\right]\nabla_{j}\delta n_{\xi}(\vec{r}),$ (S16)
from where we identify the magnon diffusion tensor element as
$D_{\xi,ij}=\frac{1}{I_{0}(2\pi)^{2}}\int d^{2}kn_{\xi,k}^{0}\varepsilon_{\xi,\vec{k}}\tau_{\xi,\vec{k}}v_{i,\xi\vec{k}}v_{j,\xi\vec{k}},$ (S17)
and thus
$j_{i,\xi}=-\sum_{j}D_{\xi,ij}\nabla_{j}\delta n_{\xi}(\vec{r}).$ (S18)
Given the above considerations, we can now write a diffusion equation for each magnon branch as
$\partial_{t}\delta n_{\alpha}+\nabla\cdot\vec{j}_{\alpha}$ $=-\frac{1}{\tau_{s}}(\delta n_{\alpha}+\chi_{0}h),$ (S19)
$\partial_{t}\delta n_{\beta}+\nabla\cdot\vec{j}_{\beta}$ $=-\frac{1}{\tau_{s}}(\delta n_{\beta}-\chi_{0}h),$
where $h$ is a generalized force thermodynamically conjugated to $\delta n_{\xi}$, the static spin susceptibility is indicated by $\chi_{0}$, and the associated chemical potential is $(\delta n_{\xi}/\chi_{0}\pm h)$ [47], with the different signs corresponding to the different magnon species [46]. The spin relaxation time is indicated by $\tau_{s}$. The magnon currents are then given by
$\vec{j}_{\alpha}$ $=-\overleftrightarrow{D}_{\alpha}\cdot\nabla(\delta n_{\alpha}+\chi_{0}h),$ (S20)
$\vec{j}_{\beta}$ $=-\overleftrightarrow{D}_{\beta}\cdot\nabla(\delta n_{\beta}-\chi_{0}h).$
For the equations, we have used the fact that, because each branch has a different spin angular momentum, the static susceptibility and the chemical potential need to have opposite signs [46].
II.1 Magnon bands in the Lieb lattice altermagnet
To calculate the elements of the diffusion tensor given in Eq. (S17), we need first the magnon velocities, which in turn depend on the magnon band structure. We consider a 2D altermagnet described by a Heisenberg Hamiltonian in a Lieb lattice, shown in Fig. 2 in the main text.
$\hat{\mathcal{H}}$ $=J_{1}\sum_{\langle\vec{r},\vec{r}^{\prime}\rangle}\hat{\vec{S}}_{i}\cdot\hat{\vec{S}}_{j}+\sum_{\langle\langle\vec{r},\vec{r}^{\prime}\rangle\rangle_{\pm}}(J_{2}\pm\Delta)\hat{\vec{S}}_{i}\cdot\hat{\vec{S}}_{j}$ (S21)
$+\frac{J_{\rm{EA}}}{2}\sum_{\vec{r}}\left(\hat{S}_{z,j}\right)^{2},$
where $J_{1}$ is the nearest-neighbor exchange between sublattices with opposite spins, $(J_{2}\pm\Delta)$ are the next-nearest-neighbor exchange, which depend on the direction, and $J_{\rm{EA}}$ is the on site easy-axis anisotropy. We first decompose the Hamiltonian in terms of intra- and intersublattice exchanges:
$\hat{\mathcal{H}}$ $=J_{1}\sum_{\vec{r}_{i},\{\vec{\delta}_{e}\}}\hat{\vec{S}}_{A,\vec{r}_{i}}\cdot\hat{\vec{S}}_{B,\vec{r}_{i}+\vec{\delta}_{e}}+\frac{J_{\rm{EA}}}{2}\sum_{\vec{r}_{i}}\left[\left(\hat{S}_{z,A,\vec{r}_{i}}\right)^{2}+\left(\hat{S}_{z,B,\vec{r}_{i}}\right)^{2}\right]$ (S22)
$+\sum_{\vec{r}_{i},\{\vec{\delta}_{1}\}}\left[(J_{2}+\Delta)\hat{\vec{S}}_{A,\vec{r}_{i}}\cdot\hat{\vec{S}}_{A,\vec{r}_{i}+\vec{\delta}_{1}}+(J_{2}-\Delta)\hat{\vec{S}}_{B,\vec{r}_{i}}\cdot\hat{\vec{S}}_{B,\vec{r}_{i}+\vec{\delta}_{1}}\right]$
$+\sum_{\vec{r}_{i},\{\vec{\delta}_{2}\}}\left[(J_{2}-\Delta)\hat{\vec{S}}_{A,\vec{r}_{i}}\cdot\hat{\vec{S}}_{A,\vec{r}_{i}+\vec{\delta}_{2}}+(J_{2}+\Delta)\hat{\vec{S}}_{B,\vec{r}_{i}}\cdot\hat{\vec{S}}_{B,\vec{r}_{i}+\vec{\delta}_{2}}\right].$
The sets of vectors describing the positions of nearest and next-nearest neighbors are
$\{\vec{\delta}_{e}\}$ $:\quad\vec{\delta}_{e}^{(1)}=\frac{a}{2}\vec{e}_{x}+\frac{a}{2}\vec{e}_{z},\quad\quad\vec{\delta}_{e}^{(2)}=-\vec{\delta}_{e}^{(1)},$ (S23)
$\quad\quad\vec{\delta}_{e}^{(3)}=\frac{a}{2}\vec{e}_{x}-\frac{a}{2}\vec{e}_{z},\quad\quad\vec{\delta}_{e}^{(4)}=-\vec{\delta}_{e}^{(3)}.$
$\{\vec{\delta}_{1}\}$ $:\quad\vec{\delta}_{1}^{(1)}=a\vec{e}_{x},\quad\quad\vec{\delta}_{1}^{(2)}=-\vec{\delta}_{1}^{(1)}.$
$\{\vec{\delta}_{2}\}$ $:\quad\vec{\delta}_{2}^{(1)}=a\vec{e}_{z},\quad\quad\vec{\delta}_{2}^{(2)}=-\vec{\delta}_{2}^{(1)}.$
We then use the standard Holstein-Primakoff method, by first writting the spin operators in terms of the bosonic creation and annhiliation operator $\hat{a}_{\vec{r}_{i}}$ and $\hat{b}_{\vec{r}_{i}}$:
$\hat{S}_{A,\vec{r}_{i}}^{(+)}=\sqrt{2s}\hat{a}_{\vec{r}_{i}},\hskip 14.22636pt\hat{S}_{A,\vec{r}_{i}}^{(-)}$ $=\sqrt{2s}\hat{a}_{\vec{r}_{i}}^{\dagger},\hskip 14.22636pt\hat{S}_{A,\vec{r}_{i}}^{(z)}=s-\hat{a}_{\vec{r}_{i}}^{\dagger}\hat{a}_{\vec{r}_{i}},$ (S24)
$\hat{S}_{B,\vec{r}_{i}}^{(+)}=\sqrt{2s}\hat{b}_{\vec{r}_{i}}^{\dagger},\hskip 14.22636pt\hat{S}_{B,\vec{r}_{i}}^{(-)}$ $=\sqrt{2s}\hat{b}_{\vec{r}_{i}},\hskip 14.22636pt\hat{S}_{B,\vec{r}_{i}}^{(z)}=-s+\hat{b}_{\vec{r}_{i}}^{\dagger}\hat{b}_{\vec{r}_{i}},$
here $\hat{a}_{\vec{r}_{i}}$ and $\hat{b}_{\vec{r}_{i}}$ We substitute the above expressions into Eq. (S22) and Fourier transform the operators:
$\hat{a}_{\vec{r}_{i}}=\frac{1}{\sqrt{N}}\sum_{\vec{k}}e^{i\vec{k}\cdot\vec{r}_{i}}\hat{a}_{\vec{k}},$ (S25)
with a similar expression for $\hat{b}$. We get the following bosonic Hamiltonian
$\frac{\hat{\mathcal{H}}}{\hbar}$ $=\sum_{\vec{k}}\left[\omega_{a}(\vec{k})\hat{a}^{\dagger}_{\vec{k}}\hat{a}_{\vec{k}}+\omega_{b}(\vec{k})\hat{b}^{\dagger}_{\vec{k}}\hat{b}_{\vec{k}}+g_{ab}(\vec{k})\left(\hat{a}_{\vec{k}}\hat{b}_{-\vec{k}}+\hat{a}^{\dagger}_{\vec{k}}\hat{b}^{\dagger}_{-\vec{k}}\right)\right],$ (S26)
where
$\hbar\omega_{a}(\vec{k})$ $=4s\left(J_{1}-2J_{2}+\frac{J_{\rm{EA}}}{2}\right)+8sJ_{2}\cos\left(\frac{a}{2}k_{z}+\frac{a}{2}k_{x}\right)\cos\left(\frac{a}{2}k_{z}-\frac{a}{2}k_{x}\right)+8s\Delta\sin\left(\frac{a}{2}k_{z}+\frac{a}{2}k_{x}\right)\sin\left(\frac{a}{2}k_{z}-\frac{a}{2}k_{x}\right),$ (S27)
$\hbar\omega_{b}(\vec{k})$ $=4s\left(J_{1}-2J_{2}+\frac{J_{\rm{EA}}}{2}\right)+8sJ_{2}\cos\left(\frac{a}{2}k_{z}+\frac{a}{2}k_{x}\right)\cos\left(\frac{a}{2}k_{z}-\frac{a}{2}k_{x}\right)-8s\Delta\sin\left(\frac{a}{2}k_{z}+\frac{a}{2}k_{x}\right)\sin\left(\frac{a}{2}k_{z}-\frac{a}{2}k_{x}\right),$
$\hbar g_{ab}(\vec{k})$ $=2sJ_{1}\left[\cos\left(\frac{a}{2}k_{z}+\frac{a}{2}k_{x}\right)+\cos\left(\frac{a}{2}k_{z}-\frac{a}{2}k_{x}\right)\right].$
An external magnetic field $H_{0}$ aligned with the anisotropy axis can be included by adding $\tilde{\gamma}H_{0}$ to $\hbar\omega_{a}(\vec{k})$ and subtracting $\tilde{\gamma}H_{0}$ from $\hbar\omega_{b}(\vec{k})$.
The Hamiltonian Eq. (S26) is diagonalized by a Bogoliubov transformation to the magnon modes $\hat{\alpha}$ and $\hat{\beta}$:
$\frac{\hat{\mathcal{H}}_{\rm{d}}}{\hbar}=\sum_{\vec{k}}\left[\omega_{\alpha}(\vec{k})\hat{\alpha}^{\dagger}_{\vec{k}}\hat{\alpha}_{\vec{k}}+\omega_{\beta}(\vec{k})\hat{\beta}^{\dagger}_{\vec{k}}\hat{\beta}_{\vec{k}}\right],$ (S28)
where the magnon frequencies are given by
$\omega_{\alpha}(\vec{k})$ $=\frac{\omega_{a}(\vec{k})-\omega_{b}(\vec{k})+\sqrt{(\omega_{a}(\vec{k})+\omega_{b}(\vec{k}))^{2}-4g_{ab}^{2}(\vec{k})}}{2},$ (S29)
$\omega_{\beta}(\vec{k})$ $=\frac{-(\omega_{a}(\vec{k})-\omega_{b}(\vec{k}))+\sqrt{(\omega_{a}(\vec{k})+\omega_{b}(\vec{k}))^{2}-4g_{ab}^{2}(\vec{k})}}{2}.$
The magnon bandgap $\Delta_{\rm{mag}}=\omega_{\alpha}(0)=\omega_{\beta}(0)$ is given explicitly by
$\Delta_{\rm{mag}}=\frac{4s}{\hbar}\sqrt{16J_{1}J_{\rm{EA}}+J_{\rm{EA}}^{2}},$ (S30)
while the magnon frequency splitting is
$\omega_{\alpha}(\vec{k})-\omega_{\beta}(\vec{k})=16s\Delta\sin\left(\frac{a}{2}k_{z}+\frac{a}{2}k_{x}\right)\sin\left(\frac{a}{2}k_{z}-\frac{a}{2}k_{x}\right).$ (S31)
II.2 Diffusion tensor
In this paper, we work in the high-temperature regime, from which $n^{0}_{\xi,\vec{k}}\approx\frac{k_{B}T}{\varepsilon_{\xi,\vec{k}}}$, and thus Eq. (S17) simplifies to
$D_{\xi,ij}=\frac{1}{V(2\pi)^{2}}\int d^{2}k\tau_{\xi,\vec{k}}v_{i,\xi\vec{k}}v_{j,\xi\vec{k}},$ (S32)
where $V=\int d^{2}k/(2\pi)^{2}$ is the volume of the Brillouin zone. We furthermore assume that $\tau_{\xi,\vec{k}}=\tau$, i.e., the magnon momentum relaxation time is independent of the momentum and of the magnon species. Such an assumption is justified in the high-temperature regime, in which most of the magnon relaxation is to lattice phonons. At low temperatures, other processes that are neglected in our description, such as relaxation due to magnon-magnon scattering, can become relevant [41].
With the magnon bands we can evaluate the magnon velocities, and then calculate the magnon diffusion tensor in Eq. (S17). Even though the magnon velocities have convoluted forms, we can infer dome underlying symmetry properties that are a consequence of the symmetries of the Heisenberg Hamiltonian of Eq. (4). Those a better analyzed in a coordinate system rotated by $\pi/4$, such that
$\frac{\sqrt{2}k_{z}}{2}+\frac{\sqrt{2}k_{x}}{2}$ $\rightarrow k_{x},$ (S33)
$\frac{\sqrt{2}k_{z}}{2}-\frac{\sqrt{2}k_{x}}{2}$ $\rightarrow k_{z}.$
From Eqs. (S27) and (S29) we can then infer that the magnon bands are symmetric $\omega_{\xi}(\vec{k})=\omega_{\xi}(-\vec{k})$, and related by a $\pi/2$ rotation in $k$-space:
$\omega_{\alpha}(\vec{k})=\omega_{\beta}(\mathcal{R}_{\pi/2}\vec{k}),$ (S34)
corresponding to the transformation $k_{x}\rightarrow k_{z}$, and $k_{z}\rightarrow-k_{z}$. We then identify the nodal lines
$\omega_{\alpha}(k,0)$ $=\omega_{\beta}(k,0),$ (S35)
$\omega_{\alpha}(0,k)$ $=\omega_{\beta}(0,k).$
We can then quickly infer that $v_{x,\alpha}=v_{z,\beta}$ and $v_{z,\alpha}=-v_{x,\beta}$, and thus $D_{\alpha,xy}=-D_{\beta,xy}=D_{2}$.
From the explicit form of the magnon frequencies, we further obtain that
$v_{x,\alpha}$ $=\mathbb{A}\cos(a^{\prime}k_{x})\sin(a^{\prime}k_{z})+\sin(a^{\prime}k_{x})\mathbb{F}^{(1)}(k_{x},k_{z})-\sin(a^{\prime}k_{x})\cos(a^{\prime}k_{z})\mathbb{F}^{(2)}(k_{x},k_{z}),$ (S36)
$v_{x,\beta}$ $=-\mathbb{A}\cos(a^{\prime}k_{x})\sin(a^{\prime}k_{z})+\sin(a^{\prime}k_{x})\mathbb{F}^{(1)}(k_{x},k_{z})-\sin(a^{\prime}k_{x})\cos(a^{\prime}k_{z})\mathbb{F}^{(2)}(k_{x},k_{z}),$
$v_{z,\alpha}$ $=\mathbb{A}\cos(a^{\prime}k_{z})\sin(a^{\prime}k_{x})+\sin(a^{\prime}k_{z})\mathbb{F}^{(1)}(k_{x},k_{z})-\sin(a^{\prime}k_{z})\cos(a^{\prime}k_{x})\mathbb{F}^{(2)}(k_{x},k_{z}),$
$v_{z,\beta}$ $=-\mathbb{A}\cos(a^{\prime}k_{z})\sin(a^{\prime}k_{x})+\sin(a^{\prime}k_{z})\mathbb{F}^{(1)}(k_{x},k_{z})-\sin(a^{\prime}k_{z})\cos(a^{\prime}k_{x})\mathbb{F}^{(2)}(k_{x},k_{z}).$
where $a^{\prime}=a/\sqrt{2}$, and the functions $\mathbb{F}^{(1,2)}$ are independent of $\Delta$ and exhibit the following properties
$\mathbb{F}^{(1,2)}(k_{x},k_{z})$ $=\mathbb{F}^{(1,2)}(k_{z},k_{x}),$ (S37)
$\mathbb{F}^{(1,2)}(-k_{x},k_{z})$ $=\mathbb{F}^{(1,2)}(k_{x},k_{z})=\mathbb{F}^{(1,2)}(k_{x},-k_{z})=\mathbb{F}^{(1,2)}(-k_{x},-k_{z}).$
For $\Delta=0$, i.e. for an antiferromagnet, considering an integration domain that is symmetric in both $k_{x}$ and $k_{z}$, one can show by decomposing the integration in Eq. (S32) in each quadrant of the $(k_{x},k_{z})$ domain, that $D_{2}=0$. Thus a non-vanishing value of $D_{2}$, within the assumptions we have adopted, is a characteristic exclusive to ALMs. The same conclusions should be valid for general d-wave ALMs.
III Altermagnetic spin diffusion response function
To obtain the spin diffusion response function, we use the magnon diffusion equations Eqs. (S19) and, where the symmetries of the ALM magnon bands, as discussed in the previous section, imply that the magnon diffusion tensors have the following form
$\overleftrightarrow{D}_{\alpha}=\begin{bmatrix}D_{1}&D_{2}\\ D_{2}&D_{1}\end{bmatrix},\quad\quad\overleftrightarrow{D}_{\beta}=\begin{bmatrix}D_{1}&-D_{2}\\ -D_{2}&D_{1}\end{bmatrix}$. (S38)
We then write a diffusion equation for $\mathcal{S}_{\parallel}=\hbar(-\delta n_{\alpha}+\delta n_{\beta})$:
$\partial_{t}\mathcal{S}_{\parallel}-\nabla\cdot\left[\overleftrightarrow{D}_{1}\cdot\nabla(\mathcal{S}_{\parallel}-\hbar\chi_{0}h)\right]+\nabla\cdot\left[\overleftrightarrow{D}_{2}\cdot\nabla\mathcal{S}_{\perp}\right]=-\frac{1}{\tau}(\mathcal{s}_{\parallel}-\hbar\chi_{0}h),$ (S39)
where $\mathcal{S}_{\perp}=\hbar(\delta n_{\alpha}+\delta n_{\beta})$ satisfies the equation
$\partial_{t}\mathcal{S}_{\perp}-\nabla\cdot\left[\overleftrightarrow{D}_{1}\cdot\nabla\mathcal{S}_{\perp}\right]+\nabla\cdot\left[\overleftrightarrow{D}_{2}\cdot\nabla(\mathcal{S}_{\parallel}-\chi_{0}h)\right]=-\frac{1}{\tau}\mathcal{S}_{\perp},$ (S40)
and we have defined
$\overleftrightarrow{D}_{1}=\begin{bmatrix}D_{1}&0\\ 0&D_{1}\end{bmatrix}\quad\quad\overleftrightarrow{D}_{2}=\begin{bmatrix}-D_{2}&0\\ 0&D_{2}\end{bmatrix}$. (S41)
Equations (S39) and (S40) are obtained by writting the diffusion equations in a basis that diagonalizes the magnon diffusion tensors $\overleftrightarrow{D}_{\alpha,\beta}$. We can solve this set of equations in the Fourier domain, which gives
$\mathcal{S}_{\parallel}(\omega,\vec{k})=\chi_{\parallel}(\omega,\vec{k})h$ (S42)
where the response function is given by Eq. (5) of the main text:
$\chi_{\parallel}(\omega,\vec{k})=\frac{\hbar\chi_{0}\left[D_{1}k^{2}+\frac{1}{\tau_{s}}-\frac{D_{2}^{2}k^{4}\cos^{2}(2\phi_{k})}{-i\omega+D_{1}k^{2}+1/\tau_{s}}\right]}{-i\omega+D_{1}k^{2}+\frac{1}{\tau_{s}}-\frac{D_{2}^{2}k^{4}\cos^{2}(2\phi_{k})}{-i\omega+D_{1}k^{2}+1/\tau_{s}}},$ (S43)
where $\phi_{k}=\rm{arctan}(k_{x}/k_{z})$.
IV Relaxation rate
With the response function Eq.  (S43), we can now calculate the QI relaxation rate with Eq. (2) of the main text:
$\Gamma[\omega]=\frac{\hbar\gamma^{2}\tilde{\gamma}^{2}}{2}{\rm{coth}}\left(\frac{\hbar\beta\omega}{2}\right)\int\frac{d^{2}\vec{k}}{(2\pi)^{2}}\mathcal{C}_{\{\Phi\}}(d,\vec{k})\chi^{\prime\prime}_{\parallel}(\omega,\vec{k}).$ (S44)
We first recall that the geometric factor is given by $\mathcal{C}_{\{\Phi\}}(d,\vec{k})=(2\pi)^{2}k^{2}e^{-2kd}\mathcal{F}_{\{\Phi\}}(\phi_{k})$, where the function $\mathcal{F}_{\{\Phi\}}(\phi_{k})$ depends only on the orientation angles, and we work in the high-temperature regime, such that:
$\Gamma[\omega]=\frac{\gamma^{2}\tilde{\gamma}^{2}}{\beta\omega}\int d\phi_{k}dkk^{3}e^{-2kd}\mathcal{F}_{\{\Phi\}}(\phi_{k})\chi^{\prime\prime}_{\parallel}(\omega,\vec{k}).$ (S45)
The imaginary part of the susceptibility is given explicitly by
$\chi_{\parallel}^{\prime\prime}=\frac{\hbar\chi_{0}\omega(D_{1}k^{2}+\frac{1}{\tau_{s}})\left[\omega^{2}+(D_{1}k^{2}+\frac{1}{\tau_{s}})^{2}\right]\left[\omega^{2}+(D_{1}k^{2}+\frac{1}{\tau_{s}})^{2}-D_{2}^{2}k^{4}\cos^{2}(2\phi_{k})\right]}{\omega^{2}\left[\omega^{2}+(D_{1}k^{2}+\frac{1}{\tau_{s}})^{2}+D_{2}^{2}k^{4}\cos^{2}(2\phi_{k})\right]^{2}+(D_{1}k^{2}+\frac{1}{\tau_{s}})^{2}\left[\omega^{2}+(D_{1}k^{2}+\frac{1}{\tau_{s}})^{2}-D_{2}^{2}k^{4}\cos^{2}(2\phi_{k})\right]^{2}}.$ (S46)
We write this function in terms of the dimensionless quantities
$\tilde{k}=kl_{0},\quad\quad\tilde{\omega}=\omega\tau_{s}\quad\quad\tilde{D}_{2}=D_{2}/D_{1},$ (S47)
where $l_{0}=\sqrt{D_{1}\tau_{s}}$, such that
$\frac{\chi_{\parallel}^{\prime\prime}}{\hbar\chi_{0}}=\tilde{\chi}_{\parallel}^{\prime\prime}=\frac{\tilde{\omega}(\tilde{k}^{2}+1)\left[\tilde{\omega}^{2}+(\tilde{k}^{2}+1)^{2}\right]\left[\tilde{\omega}^{2}+(\tilde{k}^{2}+1)^{2}-\tilde{D}_{2}^{2}\tilde{k}^{4}\cos^{2}(2\phi_{k})\right]}{\tilde{\omega}^{2}\left[\tilde{\omega}^{2}+(\tilde{k}^{2}+1)^{2}+\tilde{D}_{2}^{2}\tilde{k}^{4}\cos^{2}(2\phi_{k})\right]^{2}+(\tilde{k}^{2}+1)^{2}\left[\tilde{\omega}^{2}+(\tilde{k}^{2}+1)^{2}-\tilde{D}_{2}^{2}\tilde{k}^{4}\cos^{2}(2\phi_{k})\right]^{2}}.$ (S48)
We can then write the relaxation rate as
$\Gamma[\omega]=\frac{\hbar\tau_{s}\chi_{0}\gamma^{2}\tilde{\gamma}^{2}}{\beta l_{0}^{4}}\left[\frac{1}{\tilde{\omega}}\int d\phi_{k}d\tilde{k}\tilde{k}^{3}e^{-2\tilde{k}\tilde{d}}\mathcal{F}_{\{\Phi\}}(\phi_{k})\tilde{\chi}^{\prime\prime}_{\parallel}(\tilde{\omega},\vec{\tilde{k}})\right]=\Gamma_{c}\tilde{\Gamma}[\omega].$ (S49)
In the above equation $\Gamma_{c}=\hbar k_{B}T\tau_{s}\chi_{0}\gamma^{2}\tilde{\gamma}^{2}/l_{0}^{4}$ is a characteristic relaxation rate and $\tilde{d}=d/l_{0}$. All the plots shown in the main text are obtained via numerical integration of the above equation using standard python libraries.
V Spin conductivity
To estimate the characteristic relaxation rate $\Gamma_{c}$, we need the static spin susceptibility $\chi_{0}$, which is given by the ratio $D_{1}/\sigma$ between the diffusion coefficient $D_{1}$ and the spin conductivity $\sigma$. While a full evaluation of $\sigma$ is out of scope of this paper, we can estimate it by using the procedure in [30] for antiferromagnets. From Einstein’s relation $\sigma=D_{1}\partial\rho/\partial\mu$, where $\mu$ is the chemical potential and $\rho$ is the non-equilibrium magnon spin density, which can be written as
$\rho=\int\frac{d^{2}k}{(2\pi)^{2}}\left[\frac{1}{e^{\beta\hbar\omega_{\alpha,k}-\mu}-1}-\frac{1}{e^{\beta\hbar\omega_{\beta,k}+\mu}-1}\right].$ (S50)
For the purposes of a simple estimate, we consider a low-momentum expansion of the magnon frequencies. We use for convenience the coordinate system of Sec. IIB
$\omega_{\alpha}$ $=-\Delta_{\rm{ALM}}k_{x}k_{z}+\sqrt{v^{2}k^{2}+\Delta^{2}},$ (S51)
$\omega_{\beta}$ $=\Delta_{\rm{ALM}}k_{x}k_{z}+\sqrt{v^{2}k^{2}+\Delta^{2}}.$
We do not consider the first contribution in the above expressions due to the ALM band splitting, such that
$\frac{\partial\rho}{\partial\mu}=\frac{(k_{B}T)^{2}}{\pi^{2}v^{3}}\left[{\rm{Li}}_{2}(e^{-\beta\Delta_{\alpha}})+{\rm{Li}}_{2}(e^{-\beta\Delta_{\beta}})\right],$ (S52)
where $\Delta_{\alpha,\beta}=\Delta\pm\mu/\beta$. Typically, the chemical potential is induced by an external field which is much weaker than the magnon gap $\Delta$, as we have estimated in the main text. We thus set $\Delta_{\alpha,\beta}=\Delta$. With Eqs. (S52),(S51) we can estimate the spin conductivity which, in units of electric conductivity, is $\sim 3.8\times 10^{6}$ S/m for the parameters used in the main text and at $200$ K.
The calculations presented here are a rough estimate of the spin conductivity used as a starting point to assess the feasibility of our scheme. A more complete description can be done within the framework of, e.g. [47], which also accommodates the ballistic spin transport regime. An extension to such a complete framework is postponed to a future work.